OpenStudy (anonymous):

#1. e^7x/x^6+1 #2. tan^-1 (3x) #3. 7sin(x)sin^-1(x) #4. arcsin^2(6x+6) #5. 5cos(6 ln (x)) #6. 3sqrt4^x+7 #7. e^3x(x^2+6^x #8. Find dy/dx in terms of x and y if ln y +y^3=6 ln x #9. Find dy/dx in terms of x and y if 5xy+9x+y=8 #10. 2 ln (sinx) f"(x) ?

OpenStudy (across):

What is this?

OpenStudy (anonymous):

this is derivatives. I posted a comment before this. I have most of the answers I wanted to go over them with someone on here.

OpenStudy (across):

Okay, so, for the first one, you have to find the derivative of\[f(x)=\frac{e^{7x}}{x^6}+1?\]

OpenStudy (anonymous):

the bottom is x^6+1

OpenStudy (anonymous):

I have to find derivatives for all 10 questions. but #10 i have to find the 2nd derivative. im doing something wrong but I cant figure out what it is.

OpenStudy (across):

\[f(x)=\frac{e^{7x}}{x^6+1}.\]Well, we need to know the quotient rule for this one. That is,\[f'(x)=\frac{(x^6+1)(e^{7x})'-(x^6+1)'(e^{7x})}{(x^6+1)^2}.\]That boils down to something like\[f'(x)=\frac{(x^6+1)(7e^{7x})-(6x)(e^{7x})}{(x^6+1)^2}.\]It's just a matter of simplifying things now...

OpenStudy (anonymous):

Im good with the simplifying. I just needed to know how to get to that part.

OpenStudy (across):

10.\[f(x)=2\ln(\sin(x)).\]This requires we perform the chain rule to find the first derivative.\[\frac{df}{dx}=\frac{d}{dx}[2\ln(\sin(x))],\]\[\frac{df}{dx}=2\frac{d}{dx}[\ln(\sin(x))],\]\[\frac{df}{dx}=2\frac{1}{\sin(x)}\frac{d}{dx}[(\sin(x))],\]\[\frac{df}{dx}=2\frac{1}{\sin(x)}(\cos(x)),\]\[\frac{df}{dx}=2\cot(x).\]Can you find the derivative of the above expression?

OpenStudy (anonymous):

-csc^2x ?

OpenStudy (across):

Very close:\[-2\cdot\csc^2(x).\]

OpenStudy (anonymous):

ok. thanx. Lets go over the others??

OpenStudy (across):

Which other did you find difficult?

OpenStudy (anonymous):

well, for # 2 i got sec^2(3^x) ??

OpenStudy (across):

Hmm, I wonder how you got that.

OpenStudy (anonymous):

the derivastive of tan is sec^2 right?

OpenStudy (across):

2.\[f(x)=\tan^{-1}(3x).\]This one follows straight from the definition of the derivative of the inverse tangent function (if you don't feel like deriving it, that is).\[\frac{df}{dx}=\frac{d}{dx}[\tan^{-1}(3x)],\]\[\frac{df}{dx}=\frac{1}{1+(3x)^2}\frac{d}{dx}[3x],\]\[\frac{df}{dx}=\frac{3}{1+9x^2}.\]I think you confused these two:\[\tan(x)\neq\tan^{-1}(x).\]

OpenStudy (across):

For the record:\[\frac{d}{dx}[\tan^{-1}(x)]=\frac{1}{1+x^2}.\]

OpenStudy (anonymous):

the answer i put in was (3) / (1+9x^2) but that is only part of the answer? right? Im not sure how to combine the two. does that make sense?

OpenStudy (across):

You mean, combine the terms in the denominator?

OpenStudy (anonymous):

i mean take 1/1+x^2 and combine it with 3/1+9x^2 . Is that what im supposed to do to get my answer?

OpenStudy (across):

No, I was merely stating that\[\frac{d}{dx}[\tan^{-1}(x)]=\frac{1}{1+x^2}.\]However,\[\frac{d}{dx}[\tan^{-1}(3x)]=\frac{3}{1+9x^2},\]by applying the chain rule (see above).

OpenStudy (anonymous):

ok. could you help me with #8 and 9? what does it mean "in terms of x and y?"

OpenStudy (across):

Questions 8 and 9 are asking you to perform implicit differentiation. I suppose you have heard of it before?

OpenStudy (across):

Say I give you an example, and you follow the train of thought to solve number 9? Suppose I give you this:\[25x^2+y^2=109.\]Finding an implicit differential of this function, assuming y is the independent variable, goes as follows:\[\frac{d}{dx}[25x^2+y^2]=\frac{d}{dx}[109].\]Evaluating as usual should give you:\[{50x+2y\frac{dy}{dx}}=0.\]You then solve for dy/dx and obtain:\[\frac{dy}{dx}=-\frac{50x}{2y}=-25\frac{x}{y}.\]That's all there is to it.

OpenStudy (anonymous):

cool!

OpenStudy (across):

I meant to say "assuming y is the dependent variable." -.- Let me know when you have given those a try.

OpenStudy (anonymous):

ok...i have to jump off here for a few minutes. thanx for all your help.

OpenStudy (across):

No, you're welcome.